OpenStudy (anonymous):

How do you integrate7^2x

OpenStudy (anonymous):

\[\int b^xdx=\frac{b^x}{\ln(b)}\] is good to know in general

OpenStudy (anonymous):

i guess you also need the chain rule aka u substitution because you have \(\int 7^{2x}dx\) but that is a small challenge , make \(u=2x.\frac{1}{2}du = dx\) etc

OpenStudy (anonymous):

Could you go through the u substitution please. I have trouble with that

OpenStudy (anonymous):

\[u=2x, du = 2xdx,\frac{1}{2}du=dx\] you get \[\int7^{2x}dx=\frac{1}{2}\int 7^udu=\frac{7^u}{2\ln(7)}=\frac{7^{2x}}{2\ln(7)}\]

OpenStudy (anonymous):

sorry first line should be \(u=2x,du=2dx\)

OpenStudy (anonymous):

ok thanks a lot.

OpenStudy (anonymous):

yw